An angle is greater than 45^o . Is its complementary angle greater than 45^o or equal to 45^o or less than 45^o ?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the complementary angles be p and q,
We know that, sum of measures of complementary angle pair is 90 . Then,
= p + q = 90^o
It is given in the question that p > 45^o
Adding q on both the sides,
= p + q > 45^o+ q
= 90^o> 45^o+ q
= 90^o–45^o> q
= q < 45^o

Thanks